the difference between 2 numbers is 24. if the larger number is divided by the smaller number the quotient is 4 with a remainder of 3. find the numbers.
Answer by mananth(16949)

let the numbers be x & y , x>y
x-y =24..........(1)
4y+3=x
4y-x=-3..........(2)

Add (1) & (2)
3y=21
/3
y=7
plug y in (1)
we get x=31
HAPPY NEW YEAR